# Knowledge Mechanisms in Large Language Models: A Survey and Perspective

### v2

@@ 1 Introduction
As we know, Large language models (LLMs) are renowned for encapsulating extensive parametric knowledge Roberts et al. (2020); Sung et al. (2021); Cao et al. (2021a); Zhong et al. (2021); Kandpal et al. (2023); Heinzerling and Inui (2020); Petroni et al. (2019); Qiao et al. (2023); Kritharoula et al. (2023); He et al. (2024a), achieving unprecedented progress in application. However, the knowledge mechanisms in LLMs for learning, storage, utilization, and evolution still remain mysterious Gould et al. (2023a). Extensive works aim to demystify various types of knowledge in LLMs through knowledge neurons Dai et al. (2022); Chen et al. (2024a) and circuits Elhage et al. (2021); Yao et al. (2024); Zou et al. (2024), yet these efforts, scattered across various tasks, await comprehensive review and analysis.
@@
unsafety Wang et al. (2024b); Wu et al. (2023), language Tang et al. (2024); Zhao et al. (2023a)
@@
Previous interpretability surveys typically aim to investigate various methods for explaining the roles of different components within LLMs from the global and local taxonomy Ferrando et al. (2024); Zhao et al. (2024a); Luo and Specia (2024); Murdoch et al. (2019); Bereska and Gavves (2024); Vilas et al. (2024); Singh et al. (2024). In contrast, this paper focuses on knowledge in LLMs. Hence, our taxonomy, oriented from target knowledge in LLMs, reviews how knowledge is acquired, stored, utilized, and subsequently evolves. Additionally, previous taxonomy mostly explore the explainability during the inference stage (a specific period), while ignoring knowledge acquisition during the pre-training stage and evolution during the post-training stage Räuker et al. (2023); Luo et al. (2024b); Apidianaki (2023); Jiao et al. (2023); Räuker et al. (2023); Rai et al. (2024). Our taxonomy aims to explore the dynamic evolution across all periods from naivety to sophistication in both individual and group LLMs. In contrast to the most similar survey Cao et al. (2024a) that introduces knowledge life cycle, our work focuses on the underlying mechanisms at each stage.
@@ 2.2 Definition of Knowledge in LLMs
$\mathcal{F}\left(r_{k\backslash t}\right)=t$ (1)
@@ 2.3 The Architecture of LLMs
The prevailing architecture in current LLMs is the Transformer Vaswani et al. (2017). Specifically, a transformer-based LLM $\mathcal{F}$ begins with token embedding, followed by $L$ layers transformer block, and ends with token unembedding used for predicting answer tokens. Each block layer $l$ consists of Attention Heads (Attention) and Multilayer Perceptron (MLP):
@@ 2.4 Knowledge Analysis Methods
Knowledge analysis method $\mathcal{M}$ aims to reveal precise causal mechanisms from inputs to outputs Bereska and Gavves (2024). Furthermore, if components $\mathbf{C}$ of $\mathcal{F}$ accurately infer $t$ through analysis method $\mathcal{M}$, it is assumed that the knowledge $k$ is presented by $\mathbf{C}$:
@@
The elements in set $\mathbf{C}$ may be individual neurons, MLPs, attention heads, a transformer block layer, or knowledge circuit Yao et al. (2024). Inspired by Levels of Analysis Yurdusev (1993), these methods can be classified into two main categories: observation and intervention Bereska and Gavves (2024).
@@
$E$ is a evaluation metric, which can be a probe Räuker et al. (2023), logit lens nostalgebraist (2020), or a sparse representation Gao et al. (2024b). Probe is a meticulously trained classifier, and its classification performance is used to observe the relationship between model’s behavior and the output of $\mathbf{C}$ Belinkov (2022); Elazar et al. (2021); McGrath et al. (2021); Gurnee et al. (2023). Logit lens usually translate output of $\mathbf{C}$ into vocabulary tokens via token unembedding Geva et al. (2022b); Belrose et al. (2023); Pal et al. (2023); Din et al. (2024); Langedijk et al. (2023). Sparse representation disentangles the output of $\mathbf{C}$ into basic features via sparse autoencoder Sharkey et al. (2022); Cunningham et al. (2023); Lee et al. (2006) or sparse dictionary learning He et al. (2024b); Olshausen and Field (1997); Yun et al. (2021). Each basic feature represents an interpretable unit, and the output of $\mathbf{C}$ is the combination Elhage et al. (2022); Bricken et al. (2023) of these basic features.
@@ 3.1 Memorization
Knowledge memorization aims to remember and recall knowledge, e.g., specific terms (entities), grammar, facts, commonsense and concepts Allen-Zhu and Li (2023a); Yu et al. (2023a); Mahowald et al. (2023); Zhu and Li (2023); Allen-Zhu and Li (2023b, 2024); Cao et al. (2024a). We posit knowledge memorization from Modular Region and Connection Hypothesis by reviewing existing research.
@@
This modular region hypothesis simplifies knowledge representation in transformer-based models into isolated modular region, e.g., MLPs or attention heads. Knowledge is encoded via MLPs. Geva et al. (2021) posit that MLPs operate as key-value memories and each individual key vector corresponds to a specific semantic pattern or grammar. Based on the above finding, Geva et al. (2022b, a) reverse engineer the operation of the MLPs layers and find that MLPs can promote both semantic (e.g., measurement semantic including kg, percent, spread, total, yards, pounds, and hours) and syntactic (e.g., adverbs syntactic including largely, rapidly, effectively, previously, and normally) concepts in the vocabulary space. Miller and Neo (2024) find a single MLP neuron (in GPT-2 Large) capable of generating “an” or “a”. Subsequently, fact Dai et al. (2022); Meng et al. (2022) and commonsense knowledge Gupta et al. (2023) are found. Advanced language-specific neurons Tang et al. (2024), linguistic regions Zhao et al. (2023a), entropy neurons Stolfo et al. (2024), abstract conceptual Wang et al. (2024e) and unsafe Wang et al. (2024b); Wu et al. (2023) knowledge, are also observed in MLPs. In addition to MLP, knowledge is also conveyed by attention heads Geva et al. (2023); Gould et al. (2023b). Hoover et al. (2020) explain the knowledge each attention head has learned. Specifically, attention heads store evident linguistic features, positional information, and so on. Besides, fact knowledge Yu et al. (2023c); Li et al. (2023a) and bias Hoover et al. (2020) are mainly convey by attention heads. Jiang et al. (2024b) further observe that LLMs leverage self-attention to gather information through certain tokens in the contexts, which serve as clues, and use the value matrix for associative memory. Later, Zhu et al. (2024) also find that attention heads can simulate mental state and activate “Theory of Mind” (ToM) capability.
@@
Geva et al. (2023) outline the encoding of factual knowledge (e.g., “The capital of Ireland is Dublin”) through the following three steps: (1) subject (Ireland) information enrichment in MLPs, (2) the relation (capital of) propagates to the last token, (3) object (Dublin) is extracted by attention heads in later layers. This claim is supported by Li et al. (2024c). Similarly, Lv et al. (2024) conclude that task-specific attention head may move the topic entity to the final position of the residual stream, while MLPs conduct relation function. Moreover, the recent prominent knowledge circuit framework Nainani (2024); Yao et al. (2024); He et al. (2024b); Elhage et al. (2021); Marks et al. (2024) advocates leveraging a critical computational subgraph among all components to explore internal knowledge within LLM parameters. The competencies for indirect object identification and color object tasks are discovered to be embedded in specialized knowledge circuits Conmy et al. (2023); Wang et al. (2023c); Merullo et al. (2023a); Yu et al. (2024c). Lan et al. (2024) also identify number-related circuits that encode the predictive ability of Arabic numerals, number words, and months. More importantly, experimental evidence demonstrates that various types of knowledge, including linguistic, commonsense, factual, and biased information, are encapsulated in specific knowledge circuits Yao et al. (2024). Interestingly, knowledge encoded by specific circuits can rival or even surpass that of the entire LLM. This may be because knowledge circuits memorized the relevant knowledge, while noise from other components might impede the model’s performance on these tasks.
@@ 3.2 Comprehension and Application
Knowledge comprehension and application focus on demonstrating the understanding of memorized knowledge and then solving problems in new situations, e.g., reasoning Hou et al. (2023) and planning McGrath et al. (2021). Merrill et al. (2023) denote the transition from memorization to comprehension and application as grokking, and suggest that the grokking derives from two largely distinct subnetworks competition. Intuitively, only knowledge that is correctly memorized Prashanth et al. (2024) in §3.1 can be further applied to solving complex tasks. Therefore, we posit Reuse Hypothesis from two knowledge memorization perspectives.
@@
From the Modular Region Perspective, knowledge utilization reuses some regions. These regions might include a few neurons, attention heads, MLPs, or partial knowledge circuits. Generally, basic knowledge (position information, n-gram pattern, syntactic features) tends to be stored at earlier layers, while sophisticated knowledge (mental state, emotion, and abstract concept, e.g., prime number, Camelidae, and safety) is located at later layers Zhu et al. (2024); Jin et al. (2024a); Wang et al. (2024b, e); Men et al. (2024); Kobayashi et al. (2023). Therefore, neurons of earlier layers related to basic knowledge tend to be reused Kang and Choi (2023); Zhao et al. (2024a); Kandpal et al. (2023). Various math reasoning tasks also utilize the attention mechanism in initial layers to map input information to the final token positions, subsequently generating answers using a set of MLPs in later layers Stolfo et al. (2023); Hanna et al. (2023); Langedijk et al. (2023). Besides, some specific function regions are also reused. Specifically, retrieval heads Li et al. (2023a) are reused for Chain-of-Thought (CoT) reasoning and long-context tasks. These retrieval heads are found in 4 model families, 6 model scales, and 3 types of fine-tuning. Subsequently, induction heads, identified in Llama and GPT, are claimed to be responsible for in-context learning (ICL) tasks Olsson et al. (2022); Crosbie and Shutova (2024). Moreover, Lv et al. (2024) find that an attention head can map country names to capital cities. Li et al. (2023a) demonstrate that some attention heads (in Alpaca, Llama, and Vicuna) exhibit the capability for truthful answers instead of hallucinations. Tang et al. (2024) find language-specific neurons (in Llama and BLOOM) that process multiple languages, such as English, French, Mandarin, and others. Zhao et al. (2023a) further reveal linguistic regions (in Llama) correspond to linguistic competence, which is the cornerstone for performing various tasks. From the Connection Perspective, knowledge utilization shares partial knowledge circuits. For instance, similar tasks share subgraphs (computational circuits) with analogous roles Lan et al. (2024). Besides, knowledge circuits (in GPT2) are reused to solve a seemingly different task, e.g., indirect object identification and colored objects tasks Merullo et al. (2023a). Wang et al. (2024a) further observe that two-hop composition reasoning tasks reuse the knowledge circuits from the first hop. Yao et al. (2024) also believe that this reuse phenomenon exists in factual recall and multi-hop reasoning. Specifically, sub-circuits are reused in similar factual knowledge, such as tasks related to “city_in_country”, “name_birth_place”, and “country_language”. Besides, Dutta et al. (2024) demystify LLMs how to perform CoT reasoning, i.e., Llama facilitates CoT tasks via multiple parallel circuits enjoying significant intersection.
@@ 3.3 Creation
Knowledge creation Runco and Jaeger (2012); Sternberg (2006); Gaut (2010) emphasizes the capacity and process of forming novel and valuable things, rather than the existing ones (i.e., LLMs have seen) discussed in §3.1 and §3.2. The creations encompass two levels: 1) LLMs create new terms following the current world’s principles comprehended by LLMs, such as new proteins Shin et al. (2021); Madani et al. (2023), molecules Bagal et al. (2022); Fang et al. (2023); Edwards et al. (2022), code DeLorenzo et al. (2024), video Kondratyuk et al. (2023), models Zheng et al. (2024), names for people and companies, written stories Gómez-Rodríguez and Williams (2023); Buz et al. (2024), synthetic data Stenger et al. (2024); Mumuni et al. (2024); Abufadda and Mansour (2021), etc. These novel items operate according to the existing rules, e.g., law of conservation of energy, reasoning logic Wang et al. (2024a), or principles of probability theory. 2) LLMs may generate new rules, such as mathematical theorems, and the resulting terms will operate according to the new rules. We posit that the knowledge creation of LLMs may derive from the Extrapolation Hypothesis.
@@
However, from a statistical perspective, the intricate connections and activations between neurons, though not infinite, resist exhaustive enumeration. In terms of value, not all creations are valuable. Obtaining something valuable with an exceedingly low probability is impractical, as even a monkey could theoretically print Shakespeare’s works. How do LLMs ensure the probability of generating valuable creations? What are the mechanisms underlying the novelty and value of creation? A prevalent conjecture posits that novelty is generated through the random walk Sæbø and Brovold (2024). However, intuitively, current LLMs themselves seem unable to evaluate the value of creations due to architectural limitations Chakrabarty et al. (2024). Because, once the next token is generated, there is no intrinsic mechanism for accepting or rejecting the creations. This hinders the evaluation of the usefulness and value of proposed novelties, as humans do, by bending, blending, or breaking biases Sæbø and Brovold (2024). Some works assume that each token is indeed valuable and meets long-term expectations. However, the well-known hallucination problem Xu et al. (2024d) of LLMs refutes this assumption. Besides, the transformer architecture struggles with long context Li et al. (2024a), despite the existence of many variants for addressing this issue Huang et al. (2023c); Liu et al. (2024b); Zhang et al. (2024b); Fountas et al. (2024). More importantly, MLPs of Transformer may also work contrary to creativity, i.e., the increased attentions narrow the conditional distribution for token prediction Sæbø and Brovold (2024).

### v3

@@ 1 Introduction
As we know, Large language models (LLMs) are renowned for encapsulating extensive parametric knowledge Roberts et al. (2020); Sung et al. (2021); Cao et al. (2021a); Zhong et al. (2021); Kandpal et al. (2023); Heinzerling and Inui (2020); Petroni et al. (2019); Qiao et al. (2023); Kritharoula et al. (2023); He et al. (2024a), achieving unprecedented progress in application. However, the knowledge mechanisms in LLMs for learning, storage, utilization, and evolution still remain mysterious Phillips et al. (2021); Gould et al. (2023a). Extensive works aim to demystify various types of knowledge in LLMs through knowledge neurons Dai et al. (2022); Chen et al. (2024a) and circuits Elhage et al. (2021); Yao et al. (2024); Zou et al. (2024), yet these efforts, scattered across various tasks, await comprehensive review and analysis.
@@
unsafety Wang et al. (2024b); Wu et al. (2023a), language Tang et al. (2024); Zhao et al. (2023a)
@@
Previous interpretability surveys typically aim to investigate various methods for explaining the roles of different components within LLMs from the global and local taxonomy Ferrando et al. (2024); Zhao et al. (2024a); Luo and Specia (2024); Murdoch et al. (2019); Rai et al. (2024a); Bereska and Gavves (2024); Vilas et al. (2024); Singh et al. (2024). In contrast, this paper focuses on knowledge in LLMs. Hence, our taxonomy, oriented from target knowledge in LLMs, reviews how knowledge is acquired, stored, utilized, and subsequently evolves. Additionally, previous taxonomy mostly explore the explainability during the inference stage (a specific period), while ignoring knowledge acquisition during the pre-training stage and evolution during the post-training stage Räuker et al. (2023); Luo et al. (2024b); Apidianaki (2023); Jiao et al. (2023); Räuker et al. (2023); Rai et al. (2024b). Our taxonomy aims to explore the dynamic evolution across all periods from naivety to sophistication in both individual and group LLMs. In contrast to the most similar survey Cao et al. (2024a) that introduces knowledge life cycle, our work focuses on the underlying mechanisms at each stage.
@@ 2.2 Definition of Knowledge in LLMs
$t=\mathcal{F}\left(r_{k\backslash t}\right)$ (1)
@@ 2.3 The Architecture of LLMs
The prevailing architecture in current LLMs is the Transformer Vaswani et al. (2017). Specifically, a transformer-based LLM $\mathcal{F}$ begins with a token embedding, followed by $L$ layers transformer block, and ends with token unembedding used for predicting answer tokens. Each transformer block layer $l$ consists of Attention Heads (Attention) and Multilayer Perceptron (MLP):
@@ 2.4 Knowledge Analysis Methods
Knowledge analysis method $\mathcal{M}$ aims to interpret how LLMs work inside and reveal precise causal connections between specific components and outputs Bereska and Gavves (2024). Furthermore, if components $\mathbf{C}$ of $\mathcal{F}$ accurately infer $t$ through analysis method $\mathcal{M}$, it is assumed that the knowledge $k$ is presented by $\mathbf{C}$:
@@
The elements in set $\mathbf{C}$ may be individual neurons, MLPs, attention heads, a transformer block layer, or knowledge circuit Yao et al. (2024). These methods are divided into two categories: observation and intervention Bereska and Gavves (2024).
@@
$E$ is a evaluation metric, which can be a probe Räuker et al. (2023), logit lens nostalgebraist (2020), or a sparse representation Gao et al. (2024c). Probe is a meticulously trained classifier, and its classification performance is used to observe the relationship between model’s behavior and the output of $\mathbf{C}$ Belinkov (2022); Elazar et al. (2021); McGrath et al. (2021); Gurnee et al. (2023). Logit lens usually translate output of $\mathbf{C}$ into vocabulary tokens via token unembedding Geva et al. (2022b); Belrose et al. (2023); Pal et al. (2023); Din et al. (2024); Langedijk et al. (2023). Sparse representation maps the output of $\mathbf{C}$ into a higher-dimensional space with strong sparsity through dictionary learning He et al. (2024b); Olshausen and Field (1997); Yun et al. (2021); Karvonen et al. (2024), with sparse auto-encoder Sharkey et al. (2022); Cunningham et al. (2023); Lee et al. (2006); Gao et al. (2024a) being a prominent example. The higher-dimensional space represents independent (or monosemantic Bricken et al. (2023)) and interpretable features more easily Rai et al. (2024b). The output of $\mathbf{C}$ is the combination Elhage et al. (2022); Bricken et al. (2023) of these features.
@@ 3.1 Memorization
Knowledge memorization Schwarzschild et al. (2024); Prashanth et al. (2024) aims to remember and recall knowledge in the training corpus, e.g., specific terms (entities), grammar, facts, commonsense, concepts, etc Allen-Zhu and Li (2023a); Yu et al. (2023a); Mahowald et al. (2023); Zhu and Li (2023); Allen-Zhu and Li (2023b, 2024); Cao et al. (2024a). We posit knowledge memorization from Modular Region and Connection Hypothesis by reviewing existing research.
@@
This modular region hypothesis simplifies knowledge representation in transformer-based models into isolated modular region, e.g., MLPs or attention heads. Knowledge is encoded via MLPs. Geva et al. (2021) posit that MLPs operate as key-value memories and each individual key vector corresponds to a specific semantic pattern or grammar. Based on the above finding, Geva et al. (2022b, a) reverse engineer the operation of the MLPs layers and find that MLPs can promote both semantic (e.g., measurement semantic including kg, percent, spread, total, yards, pounds, and hours) and syntactic (e.g., adverbs syntactic including largely, rapidly, effectively, previously, and normally) concepts in the vocabulary space. Miller and Neo (2024) find a single MLP neuron (in GPT-2 Large) capable of generating “an” or “a”. Subsequently, fact Dai et al. (2022); Meng et al. (2022) and commonsense knowledge Gupta et al. (2023) are found. Advanced language-specific neurons Tang et al. (2024), linguistic regions Zhao et al. (2023a), entropy neurons Stolfo et al. (2024), abstract conceptual Wang et al. (2024e) and unsafe Wang et al. (2024b); Wu et al. (2023a) knowledge, are also observed in MLPs. In addition to MLP, knowledge is also conveyed by attention heads Geva et al. (2023); Gould et al. (2023b). Hoover et al. (2020) explain the knowledge each attention head has learned. Specifically, attention heads store evident linguistic features, positional information, and so on. Besides, fact knowledge Yu et al. (2023c); Li et al. (2023a) and bias Hoover et al. (2020) are mainly convey by attention heads. Jiang et al. (2024b) further observe that LLMs leverage self-attention to gather information through certain tokens in the contexts, which serve as clues, and use the value matrix for associative memory. Later, Zhu et al. (2024) also find that attention heads can simulate mental state and activate “Theory of Mind” (ToM) capability.
@@
Geva et al. (2023) outline the encoding of factual knowledge (e.g., “The capital of Ireland is Dublin”) through the following three steps: (1) subject (Ireland) information enrichment in MLPs, (2) the relation (capital of) propagates to the last token, (3) object (Dublin) is extracted by attention heads in later layers. This claim is supported by Li et al. (2024d). Similarly, Lv et al. (2024) conclude that task-specific attention head may move the topic entity to the final position of the residual stream, while MLPs conduct relation function. Moreover, the recent prominent knowledge circuit framework Nainani (2024); Yao et al. (2024); He et al. (2024b); Elhage et al. (2021); Marks et al. (2024) advocates leveraging a critical computational subgraph among all components to explore internal knowledge within LLM parameters. The competencies for indirect object identification and color object tasks are discovered to be embedded in specialized knowledge circuits Conmy et al. (2023); Wang et al. (2023c); Merullo et al. (2023a); Yu et al. (2024c). Lan et al. (2024) also identify number-related circuits that encode the predictive ability of Arabic numerals, number words, and months. More importantly, experimental evidence demonstrates that various types of knowledge, including linguistic, commonsense, factual, and biased information, are encapsulated in specific knowledge circuits Yao et al. (2024). Interestingly, knowledge encoded by specific circuits can rival or even surpass that of the entire LLM. This may be because knowledge circuits memorized the relevant knowledge, while noise from other components might impede the model’s performance on these tasks.
@@ 3.2 Comprehension and Application
Knowledge comprehension and application focus on demonstrating the understanding of memorized knowledge and then solving problems in new situations, e.g., generalization on out-of-domain tasks Wang et al. (2024a), reasoning Hou et al. (2023) and planning McGrath et al. (2021). Merrill et al. (2023) denote the transition from memorization to comprehension and application as grokking, and suggest that the grokking derives from two largely distinct subnetworks competition. Intuitively, only knowledge that is correctly memorized Prashanth et al. (2024) in §3.1 can be further applied to solving complex tasks. Therefore, we posit the following Reuse Hypothesis from two knowledge memorization perspectives.
@@
From the Modular Region Perspective, knowledge utilization reuses some regions. These regions might include a few neurons, attention heads, MLPs, a transformer layer, or partial knowledge circuits. Generally, basic knowledge (position information, n-gram pattern, syntactic features) tends to be stored at earlier layers, while sophisticated knowledge (mental state, emotion, and abstract concept, e.g., prime number, Camelidae, and safety) is located at later layers Zhu et al. (2024); Jin et al. (2024a); Wang et al. (2024b, e); Men et al. (2024); Kobayashi et al. (2023). Therefore, neurons of earlier layers related to basic knowledge tend to be reused Kang and Choi (2023); Zhao et al. (2024a); Kandpal et al. (2023). Various math reasoning tasks also utilize the attention mechanism in initial layers to map input information to the final token positions, subsequently generating answers using a set of MLPs in later layers Stolfo et al. (2023); Hanna et al. (2023); Langedijk et al. (2023). Besides, some specific function regions are also reused. Specifically, retrieval heads Li et al. (2023a) are reused for Chain-of-Thought (CoT) reasoning and long-context tasks. These retrieval heads are found in 4 model families, 6 model scales, and 3 types of fine-tuning. Subsequently, induction heads, identified in Llama and GPT, are claimed to be reused for in-context learning (ICL) tasks Olsson et al. (2022); Crosbie and Shutova (2024). Attention heads can map country names to their capitals in capital city-related tasks Lv et al. (2024). Language-specific neurons (in Llama and BLOOM) are responsible for multiple language related tasks, such as English, French, Mandarin, and others Tang et al. (2024). Zhao et al. (2023a) further reveal linguistic regions (in Llama) correspond to linguistic competence, which is the cornerstone for performing various tasks. Later, function regions related to the process of math reasoning are also discovered in LLMs. For instance, the last layer of GPT-2 (trained from scratch) has been observed to exhibit mathematical reasoning abilities across various math questions Ye et al. (2024). From the Connection Perspective, knowledge utilization shares partial knowledge circuits. For instance, similar tasks share subgraphs (computational circuits) with analogous roles Lan et al. (2024). Besides, knowledge circuits (in GPT2) are reused to solve a seemingly different task, e.g., indirect object identification and colored objects tasks Merullo et al. (2023a). Wang et al. (2024a) further observe that two-hop composition reasoning tasks reuse the knowledge circuits from the first hop. Yao et al. (2024) also believe that this reuse phenomenon exists in factual recall and multi-hop reasoning. Specifically, sub-circuits are reused in similar factual knowledge, such as tasks related to “city_in_country”, “name_birth_place”, and “country_language”. Besides, Dutta et al. (2024) demystify LLMs how to perform CoT reasoning, i.e., Llama facilitates CoT tasks via multiple parallel circuits enjoying significant intersection.
@@ 3.3 Creation
Knowledge creation Runco and Jaeger (2012); Sternberg (2006) emphasizes the capacity and process of forming novel and valuable things, rather than the existing ones (i.e., LLMs have seen) discussed in §3.1 and §3.2. The creations encompass two levels: 1) LLMs create new terms following the current world’s principles comprehended by LLMs, such as new proteins Shin et al. (2021), molecules Bagal et al. (2022); Fang et al. (2023); Edwards et al. (2022), code DeLorenzo et al. (2024), video Kondratyuk et al. (2023), models Zheng et al. (2024), names for people and companies, written stories Pépin et al. (2024); Gómez-Rodríguez and Williams (2023); Buz et al. (2024), synthetic data Stenger et al. (2024); Mumuni et al. (2024); Abufadda and Mansour (2021), etc. These novel items operate according to the existing rules, e.g., law of conservation of energy, reasoning logic Wang et al. (2024a), or principles of probability theory. 2) LLMs may generate new rules, such as mathematical theorems, and the resulting terms will operate according to the new rules. We posit that the knowledge creation of LLMs may derive from the Extrapolation Hypothesis.
@@
However, from statistical perspective, the intricate connections and activations between neurons, though not infinite, resist exhaustive enumeration. In terms of value, not all creations are valuable. Obtaining something valuable with an exceedingly low probability is impractical, as even a monkey could theoretically print Shakespeare’s works. How do LLMs ensure the probability of generating valuable creations? What are the mechanisms underlying the novelty and value of creation? A prevalent conjecture posits that novelty is generated through the random walk Sæbø and Brovold (2024). However, intuitively, current LLMs themselves seem unable to evaluate the value of creations due to architectural limitations Chakrabarty et al. (2024). Because, once the next token is generated, there is no intrinsic mechanism for accepting or rejecting the creations. This hinders the evaluation of the usefulness and value of proposed novelties, as humans do, by bending, blending, or breaking biases Sæbø and Brovold (2024). Some works assume that each token is indeed valuable and meets long-term expectations. However, the well-known hallucination problem Xu et al. (2024d) of LLMs refutes this assumption. Besides, the transformer architecture struggles with long context Li et al. (2024b), despite the existence of many variants for addressing this issue Huang et al. (2023c); Liu et al. (2024b). More importantly, MLPs of Transformer may also work contrary to creativity, i.e., the increased attentions narrow the conditional distribution for token prediction Sæbø and Brovold (2024).